OpenStudy (anonymous):

Solve negative two-thirds times a plus two-fifths equals three-fourths for a. a equals negative seven-fortieths a equals negative seven-tenths a equals negative twenty-one fortieths a equals negative twenty-one tenths

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

-2/3a + 2/5 = 3/4 --- subtract 2/5 from both sides -2/3a = 3/4 - 2/5 -- common denominator is 20 -2/3a = 15/20 - 8/20 -2/3a = 7/20 --- divide by -2/3 a = (7/20) / (-2/3) a = 7/20 * -3/2 a = - 21/40